# Path-integral Monte Carlo estimator for the dipole polarizability of quantum plasma

### v1

@@ Abstract
We present a path-integral Monte Carlo estimator for calculating the dipole polarizability of interacting Coulomb plasma in the long-wavelength limit, i.e., the optical region. Unlike the conventional dynamic structure factor in reciprocal space, our approach is based on the real-space dipole autocorrelation function and is suited for long wavelengths and small cell sizes, including finite clusters. The simulation of thermal equilibrium in imaginary time has exact Coulomb interactions and Boltzmann quantum statistics. For reference, we demonstrate analytic continuation of the Drude model into the imaginary time and Matsubara series, showing perfect agreement with our data within ranges of finite temperatures and densities. Method parameters, such as the finite time-step and finite-size effects prove only modestly significant. Our method, here carefully validated against an exactly solvable reference, remains amenable to more interesting domains in higher-order optical response, quantum confinements and quantum statistical effects, and applications in plasmonics, heterogeneous plasmas and nonlinear optics, such as epsilon-near-zero materials.
@@ I Introduction
The Drude model, introduced at the beginning of the last century by Paul Drude as an attempt at a comprehensive electron theory of metals [1, 2], remains nowadays ubiquitous in describing the optical response of metals, as well as optical materials with metal-like behaviours, such as epsilon-near-zero (ENZ) media [3]. The Drude parameters have been fitted to experiments [4] and hydrodynamic models of various kinds typically recover the Drude model or something close to it [5]. One of the advantages of the Drude model, beside its simplicity, resides in its high effectiveness in describing the optical response in a broad range of frequencies, from the near IR to the UV regime. This allows to characterize the behaviour of different photonic materials, such as metals [6, 7], plasmonic circuits [8], ENZ materials [9], as well as more complicated photonics platforms, such as metasurfaces [10], with an effective Drude model.
Thus, there is great intrigue towards the recovery of the effective response properties from first principles. In spirit, the Drude model describes the dynamics of non-interacting free electrons, which occasionally scatter from ions. When the scattering is rare, this model becomes the homogeneous electron gas (HEG), or jellium, whose quantum mechanical linear response function is the Lindhard function $\chi_{0}(\mathbf{q},\omega))$ [11], where $\mathbf{q}$ and $\omega$ treat, respectively, the momentum and the frequency of the incoming field. Not surprisingly, the Lindhard function, despite begin famously hard to evaluate generally, recovers the Drude model upon taking approximations of long wavelength, high temperature and low scattering [7, 6].
The Lindhard response can be expressed more generally as [12, 13]
$\chi(\mathbf{q},\omega)=\frac{\chi_{0}(\mathbf{q},\omega)}{1+v_{q}(1-G(\mathbf{q},\omega))\chi_{0}(\mathbf{q},\omega)},$ (1)
@@
where the fraction treats dielectric screening, $v_{q}=e^{2}/q^{2}\varepsilon_{0}$ is the reciprocal Coulomb potential (in atomic units) and $G(\mathbf{q},w)$ is the so-called local field correction, basically treating effects due to Fermion exchange and dynamic correlations. For decades, numerous works (e.g. Refs. [12, 14, 13]) have discussed and featured either static $G(\mathbf{q},\omega=0)$ or dynamic $G(\mathbf{q},\omega)$ in pursuit of the non-trivial response within the so-called regime of warm, dense matter (WDM).
@@
Recently, significant progress has been made to calculate the dielectric response of the WDM with the path-integral Monte Carlo method (PIMC) [15, 16, 17, 18, 19, 13, 20, 21]. The PIMC approach can treat accurate quantum many-body interactions at non-trivial temperatures, and it is in principle only subject to numerical challenges due to the Fermion sign problem (FSP) [22, 23]. Within this framework, moreover, the dielectric response is directly connected to the dynamic structure factor through the intermediate scattering function [11]. The latter can be straightforwardly estimated from the PIMC walker as the density-density correlation function in the imaginary time. This function can be then analytically continued in the complex plane to recover the full dynamic response of the system. The inverse transformation back to the real time, however, faces serious challenges with numerical noise, prevalent in Monte Carlo simulation, and thus requires special methods [24, 25, 20].
Within this framework, Dornheim et al. have comprehensively studied the first [13] and higher order [17, 18] response properties of the HEG. An apparent limitation of their approach is that the allowed values for the wavenumber are limited by $q=2\pi n/L$, where $n\geq 1$ and $L$ is the (cubic) simulation cell size. Therefore, the typical optical wavelengths with $q\ll L^{-1}$, thus requiring large $L$, are hard to accomplish because of the algorithmic scaling and performance loss attributed to the high particle count $N$. Indeed, the literature focuses on characterizing inelastic neutron scattering or X-ray Thompson scattering spectra, where the momentum transfers are significant [13] compared to that of photons.
@@
Here, we explore a complementary PIMC approach to regain focus on the long, optical wavelengths. Namely, we demonstrate computation of the conventional optical response based on the complex dipole polarizability, or the dipole-dipole correlation function. This approach, which evades the difficulty of low $q$, has been studied before in finite systems [26, 27, 28, 29] but not in periodic ones. Granted, formal problems arise that are more naturally addressed in the Fourier space: The long-wavelength limit is only quasi-accurate, since we treat optical fields that are locally constant inside a finite box, i.e. $\lambda\gg L$, and formally expand that by summing over infinite periodic images. In practice, there is no field in the simulation, and thus, the approximation remains justified in all parameter ranges in the intended scope: for example, the box size $L$ required for $N=16$ particles to simulate the metallic density of gold is over $300$ smaller than an ultraviolet wavelength $\lambda=200$ nm.
@@
However, it is not obvious, a priori, how a PIMC estimator for the dipole polarizability must be laid out for periodic systems or how it responds to finite-size effects or converge parameters, such as the finite time-step. Therefore, it is the aim of this work to establish these matters in a controllable setting: We first discuss analytical correspondence between the Drude model and the PIMC polarizability, evaluated in the imaginary time. For convenience and consistency with the Drude model, we use distinguishable particles, i.e. the Boltzmann statistics, to evade the Fermion sign problem—for now. We then survey various methodological and physical parameters to assert that the correspondence holds, when the Drude scattering rate is low. Indeed, we show that biases are statistically non-distinguishable within $2\%$ accuracy; the PIMC polarizability matches the Drude model. Formally, this is well anticipated, but practically the lack of biases is a signal of a robust simulation method. The assessments of performance help us avoid pitfalls, establish good practices and assert validity for continuation studies.
@@ II Theory
In this section, we first overview the linear optical response of quantum particles in the long-wavelength limit. We then consider analytic continuation of the Drude model into imaginary domain, where it meets the dipole moment autocorrelation function accessible to the PIMC method. Throughout this work, we use the Hartree atomic units: $\hbar=1/(4\pi\epsilon_{0})=m_{e}=1$.
@@ II.1 Electric field response
The first-order local polarization $\mathbf{P}(\mathbf{r},\omega)$ of charged plasma responding to an incident electric field $\mathbf{F}(\mathbf{r^{\prime}},\omega)$ can be written as
@@
$\mathbf{P}(\mathbf{r},\omega)=\varepsilon_{0}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\mathbf{F}(\mathbf{r^{\prime}},\omega),$ (2)
where $\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)$ is a complex linear response function, the electric susceptibility, in the local response approximation. Let us consider the long-wavelength limit, in which the field is oscillating with a wavelength $\lambda=2\pi c/\omega\gg L$, where $c$ is the speed of light and $L$ is a measure of longitudinal distance containing the interaction. Then, within $L$ we may consider a locally constant field
$\mathbf{F}(\mathbf{r^{\prime}},\omega)\rightarrow\delta(\mathbf{r^{\prime}}-\mathbf{r})\mathbf{F}(\omega),$ (3)
where $\delta$ is the Dirac delta function, inducing total mean polarization $P(\omega)\equiv\varepsilon_{0}\chi(\omega)\mathbf{F}(\omega)$ according to total mean susceptibility
$\chi(\omega)=\frac{1}{L^{d}}\int_{0}^{L}\chi(\mathbf{r},\mathbf{r^{\prime}},\omega)\delta(\mathbf{r}-\mathbf{r^{\prime}})d\mathbf{r^{\prime}},$ (4)
where $d$ is the dimensionality of $\mathbf{r}$. We may further consider the per-particle dipole polarizability $\alpha(\omega)$ through
$\varepsilon_{0}\chi(\omega)=n\alpha(\omega),$ (5)
where $n=N/L^{3}$ is the number density of the particles within $L$. The tensorial characters of $\chi$ and $\alpha$ are omitted here for simplicity of notation; because of the radial symmetry of the plasma, we only consider the diagonal, where $\mathbf{P}$ and $\mathbf{F}$ are parallel.
For quantized plasma represented by $N$ particles, the per-particle dipole polarizability $\alpha(\omega)$ can also be expressed using the Kubo formula [30]
$\alpha(\omega)$ $=-\frac{i}{N\hbar}\int_{0}^{\infty}\mathrm{d}te^{i\omega t}\int d\mathbf{r}\langle[\mu(\mathbf{r},t),\mu(\mathbf{r},0)]\rangle$ (6)
$\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}G^{R}_{\mu\mu}(t)$ (7)
$\equiv-G^{R}_{\mu\mu}(\omega),$ (8)
where $G^{R}$ denotes the retarded dipole-dipole correlation function and the dipole moment operator is given by
$\mu(\mathbf{r},t)=\sum_{i=1}^{N}q_{i}[\delta(\mathbf{r}-\mathbf{r_{i}}(t))-\langle\mathbf{r_{i}}\rangle],$ (9)
where $q_{i}$ is the charge and $\mathbf{r_{i}}(t)$ the time-dependent displacement of the $i$th particle, and $\langle\mathbf{r_{i}}\rangle$ is their mean displacement, a non-perturbed reference point. Clearly, when $\sum q_{i}\neq 0$ the value depends on the reference point, which is problematic in translationally invariant systems, such as plasmas. Therefore, it is more meaningful to consider the relative correlation function
$\tilde{G}^{R}_{\mu\mu}(t)\equiv G^{R}_{\mu\mu}(t)-G^{R}_{\mu\mu}(0)$ (10)
which is zero at $t=0$ but nevertheless yields
$\tilde{\alpha}(\omega)$ $\equiv-\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}\tilde{G}^{R}_{\mu\mu}(t)=-G^{R}_{\mu\mu}(\omega)=\alpha(\omega)$ (11)
for all $\omega\neq 0$, while $\alpha(\omega=0)$ remains elusive. On the other hand, it is well known that the total response of charged gas in the limits of $\omega=0$ and $q=0$ (long wavelength) is zero due to perfect screening [31]. The linear response framework considered here does not treat screening, but this should justify why $\alpha(\omega=0)$ is unimportant for the free gas.
The relative polarizability $\tilde{\alpha}(\omega)$ is analytic in the upper complex plane and can be analytically continued to imaginary time $it\rightarrow\tau$. The associated, time-ordered imaginary-time correlation function is given by
$G_{\mu\mu}(\tau)=\langle\mathcal{T}\mu(0)\mu(\tau)\rangle,$ (12)
where $\mathcal{T}$ is the time-ordering operator. The analytic continuation can be done via the spectral density function
$A(\omega)\equiv-2\mathrm{Im}[G^{R}_{\mu\mu}(\omega)]$ (13)
where the bosonic integral transformation yields
$G_{\mu\mu}(\tau)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega),$ (14)
where $G_{\mu\mu}(\tau)=G_{\mu\mu}(\tau+n\beta)$ for any integer $n$ and $\beta=1/k_{B}T$ is the inverse temperature. It is easy to see that
$\tilde{G}_{\mu\mu}(\tau)$ $\equiv G_{\mu\mu}^{R}(\tau)-G_{\mu\mu}^{R}(0)$ (15)
$=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1-\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega).$ (16)
Alternatively, we may write
$G_{\mu\mu}(i\omega_{n})=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{i\omega_{n}-\omega}A(\omega),$ (17)
where $\omega_{n}=2\pi n/\hbar\beta$ are the bosonic Matsubara frequencies.
While the imaginary-domain functions $G_{\mu\mu}(\tau)$ and $G_{\mu\mu}(i\omega_{n})$ contain the same information as $G^{R}_{\mu\mu}(\omega)$, the conversion from imaginary to real domain by inverting Eqs. (14) or (17) is a known ill-posed problem, which is accentuated by the unavoidable error bars present in Quantum Monte Carlo methods [24]. The use of existing numerical methods of analytic continuation, such as Maximum Entropy (MaxEnt) [25, 26], will be left for future consideration.
II.2 Analytic continuation of the Drude model
The ubiquitous model for considering the optical response of microscopic particles is the Drude model. The Drude susceptibility is given by
$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (18)
where $\omega_{p}^{2}=\tfrac{ne^{2}}{m\varepsilon_{0}}=\tfrac{e^{2}}{4\pi\varepsilon_{0}m}\tfrac{3}{r_{s}^{3}}$ is the plasma frequency and $\Gamma$ is a positive damping coefficient. The Drude per-particle polarizability is
$\alpha^{D}(\omega)=\frac{\chi_{e}^{D}(\omega)\varepsilon_{0}}{n},$ (19)
whose spectral density, according to Eq (13), is given by
$A^{D}(\omega)$ $=\frac{2\omega_{p}^{2}\varepsilon_{0}}{n}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}$ (20)
$=\frac{2e^{2}}{m}\frac{1}{\Gamma\omega(1+\omega^{2}/\Gamma^{2})}.$ (21)
The spectrum diverges at $\omega=0$ and its analytic continuation with Eq. (14) becomes problematic. However, the relative polarizability $\tilde{\alpha}(\omega)$ can be obtained from Eq. (16) because the divergence is controlled at $\omega=0$ by the transformation kernel.
Furthermore, an analytical expression for $G^{D}_{\mu\mu}(i\omega_{n})$, which is the correlation function in Matsubara frequencies associated to the Drude model, can be obtained by using Eq.  (20) and substituting into Eq. 17. This yields

### v2

@@ Abstract
We present a path-integral Monte Carlo estimator for calculating the dipole polarizability of interacting Coulomb plasma in the long-wavelength limit, i.e., the optical region. We present comprehensive details and method validation studies for our approach based on dipole imaginary-time autocorrelation functions. The simulation of thermal equilibrium in imaginary time has exact Coulomb interactions and Boltzmann quantum statistics. For reference, we use the Drude model as the long-wavelength limit of the Lindhard response, presenting its analytical continuation into the imaginary time and Matsubara series. We demonstrate great agreement within $1\%$ between PIMC and the reference model, indicating negligible numerical biases and physical many-body effects in metallic densities. The approach is amenable to nonlinear optical response, quantum confinements, dispersion forces, and to inform applications such as plasmonics and epsilon-near-zero materials.
@@ I Introduction
The Drude model, introduced at the beginning of the last century by Paul Drude as an attempt at a comprehensive electron theory of metals [1, 2], remains nowadays ubiquitous in describing the optical response of metals, as well as optical materials with metal-like behaviours, such as epsilon-near-zero (ENZ) media [3]. The Drude parameters have been fitted to experiments [4] and hydrodynamic models of various kinds typically recover the Drude model or something close to it [5]. One of the advantages of the Drude model, beside its simplicity, resides in its high effectiveness in describing the optical response in a broad range of frequencies, from the near IR to the UV regime. This allows to characterize the behaviour of different photonic materials, such as metals [6, 7], plasmonic circuits [8], ENZ materials [9], as well as more complicated photonics platforms, such as metasurfaces [10], with an effective Drude model. Yet, the Drude model remains phenomenological and unable to predict all subtleties of quantum mechanics from first principles.
In spirit, the Drude model describes the dynamics of non-interacting free electrons, which occasionally scatter from ions. When the scattering is modest, this model becomes the homogeneous electron gas (HEG), or jellium. Perhaps the most established framework for treating the dielectric response of HEG is the Lindhard theory, which builds upon the Lindhard function $\chi_{0}(\mathbf{q},\omega)$ [11], where $\mathbf{q}$ and $\omega$ are the respective wave number and frequency of an incoming perturbation. While somewhat tedious to evaluate numerically, $\chi_{0}$ recovers many physical insights in the right limits. In particular, the Drude model is recovered by taking $\chi_{0}$ to the limits of long wavelength and low scattering [7, 6].
The Lindhard function cannot also describe the full many-body interactions of HEG, unless the Lindhard response is expressed more generally, like [12, 11]
$\chi(\mathbf{q},\omega)$ $=-v_{q}\chi_{nn}(\mathbf{q},\omega)$ (1)
$\chi_{nn}(\mathbf{q},\omega)$ $=\frac{\chi_{0}(\mathbf{q},\omega)}{1+v_{q}G(\mathbf{q},\omega)\chi_{0}(\mathbf{q},\omega)},$ (2)
@@
where $\chi$ denotes the complex dielectric susceptibility, $\chi_{nn}$ is the screened density–density susceptibility, $v_{q}=e^{2}/|\mathbf{q}|^{2}\varepsilon_{0}$ is the reciprocal Coulomb potential (in SI units) and $G(\mathbf{q},\omega)$ is the so-called local field correction treating many-body effects due to exchange and correlation. When $G=0$, Eq. (2) reduces to $\chi_{0}(\mathbf{q},\omega)$ and amounts to the random phase approximation. For decades, numerous works (e.g. Refs. [12, 13, 14, 15]) have discussed and featured either static $G(\mathbf{q},\omega=0)$ or dynamic $G(\mathbf{q},\omega)$ in pursuit of the non-trivial response within the so-called regime of warm, dense matter (WDM).
@@
Recently, significant progress has been made to calculate the density response of the WDM with the path-integral Monte Carlo method (PIMC) [16, 17, 18, 19, 20, 15, 21, 22]. The PIMC approach can treat accurate quantum many-body interactions at finite temperatures, and it is in principle only subject to numerical challenges due to the Fermion sign problem (FSP) [23, 24]. Within this framework, moreover, the dielectric response is directly connected to the dynamic structure factor through the intermediate scattering function [11]. The latter can be straightforwardly estimated from the PIMC walker as the density-density correlation function in the imaginary time. This function can be then analytically continued in the complex plane to recover the full dynamic response of the system. The inverse transformation back to the real time, however, faces serious challenges with numerical noise, prevalent in Monte Carlo simulation, and thus requires special methods [25, 26, 21]. Within this framework, Dornheim et al. have comprehensively studied the first [15] and higher order [18, 19] response properties of the HEG. An apparent limitation of their approach is that the allowed values for the wavenumber are limited by $|\mathbf{q}|=2\pi n/L$, where $n\geq 1$ and $L$ is the (cubic) simulation cell size. Therefore, the typical optical wavelengths with $q\ll L^{-1}$, thus requiring large $L$, are hard to accomplish because of the algorithmic scaling and performance loss attributed to the high particle count $N$. Indeed, the literature focuses on characterizing inelastic neutron scattering or X-ray Thompson scattering spectra, where the momentum transfers are significant [15] compared to that of photons.
@@
Here, we explore a complementary PIMC approach to regain focus on the long, optical wavelengths. We, in fact, demonstrate computation of the conventional optical response based on the complex dipole polarizability, or the dipole-dipole correlation function. This approach, which evades the difficulty of low $\mathbf{q}$, has been studied before in finite systems [27, 28, 29, 30] but not in periodic ones. The long-wavelength approximation, where we treat a constant field in space, is well justified for optical region: For example, the box size $L$ required for $N=16$ particles to simulate the metallic density of gold is over $300$ times smaller than an ultraviolet wavelength $\lambda=200$ nm. Treating this naïvely in the Fourier space would require $N\sim 10^{8}$ particles in a cubic box.
@@
However, it is not obvious a priori how a PIMC estimator for the dipole polarizability must be laid out for periodic systems or how it responds to finite-size effects or converge parameters, such as the finite time-step. For these reasons, this work aims at establishing these matters in a controllable setting. To do so, we first discuss analytical correspondence between the Drude model and the PIMC polarizability, evaluated in the imaginary time. For convenience and consistency with the Drude model, we use distinguishable particles, i.e., Boltzmann statistics, to evade the Fermion sign problem, leaving a comprehensive investigation of its effects on our model to future works. We then survey various methodological and physical parameters to assert that the correspondence holds, when the Drude scattering rate is low. Indeed, we show that biases are statistically non-distinguishable within $1\%$ relative accuracy; the PIMC polarizability matches the Drude model at metallic densities. While the result is not overly interesting in this regime, it asserts that the numerical method is robust and accurate for numerous continuation prospects.
@@ II Theory
In this section, we first overview the linear optical response of quantum particles in the long-wavelength limit. We then discuss estimation of the dipole–dipole correlation function from imaginary-time PIMC simulation. Finally, we consider the Drude model as a non-interacting reference theory and its analytic continuation into the imaginary domain. The results of this section are presented in SI units, but the reader should keep in mind that later results will be given in in Hartree atomic units, i.e., $\hbar=4\pi\varepsilon_{0}=m_{e}=1$, for convenience.
@@ II.1 Electric field response
Let us consider a quantum mechanical plasma that is perturbed by a weak external electric field $\mathbf{E}(\mathbf{r^{\prime}},\omega)$ such that the Hamiltonian becomes
$H\rightarrow H_{0}+H_{E}(\omega),$ (3)
where $H_{0}$ is the non-perturbed Hamiltonian and $H_{E}(\omega)$ is a field-induced energy shift, integrated over volume $V$, is
$H_{E}(\omega)$ $=\int_{V}d\mathbf{r}U_{E}(\mathbf{r})$ (4)
$U_{E}(\mathbf{r})$ $=\mathbf{P}(\mathbf{r},\omega)\cdot\mathbf{E}(\mathbf{r},\omega),$ (5)
where the induced electric polarization is prescribed to be linear with the (weak) field:
$\mathbf{P}(\mathbf{r},\omega)\cong\varepsilon_{0}\int_{V}d\mathbf{r}^{\prime}\chi(\mathbf{r},\mathbf{r}^{\prime},\omega)\mathbf{E}(\mathbf{r^{\prime}},\omega),$ (6)
@@
where $\varepsilon_{0}$ the vacuum permittivity and $\chi(\mathbf{r},\mathbf{r}^{\prime},\omega)$ is the dielectric susceptibility tensor.
By assuming translation invariance of plasma and regularity of $\mathbf{E}(\mathbf{r})$, we may write $\chi(\mathbf{r},\mathbf{r}^{\prime},\omega)=\chi(\mathbf{r}-\mathbf{r}^{\prime},\omega)$ and consider its Fourier transform $\chi(\mathbf{q},\omega)$ like in Eq. (1). On the other hand, we can make a long-wavelength approximation in the locality $L\sim V^{-3}$ such that $L\sim|\mathbf{r}-\mathbf{r^{\prime}}|\ll\lambda$, where $\lambda=2\pi/|\mathbf{q}|$ is the wavelength of the locally constant field $\mathbf{E}(\mathbf{r}-\mathbf{r}^{\prime},\omega)\cong\mathbf{E}(\omega)$. Thus, we can re-express Eq. (5) as a uniform potential energy density with
$\delta U_{E}(\omega)$ $\cong\frac{1}{V}\int_{V}d(\mathbf{r}-\mathbf{r^{\prime}})\mathbf{P}(\mathbf{r}-\mathbf{r^{\prime}},\omega)\cdot\mathbf{E}(\omega)$ (7)
$=\varepsilon_{0}|\mathbf{E}(\omega)|^{2}\chi(\omega)$ (8)
where $\chi(\omega)\equiv V^{-1}\int_{V}d(\mathbf{r}-\mathbf{r^{\prime}})\chi(\mathbf{r}-\mathbf{r^{\prime}},\omega)$ is the bulk dielectric susceptibility, describing linear optical dispersion in the long-wavelength limit through the canonical relations [31]
$\varepsilon(\omega)$ $=1+\chi(\omega)$ (9)
$\mathbf{D}(\omega)$ $=\varepsilon_{0}\varepsilon(\omega)\mathbf{E}(\omega),$ (10)
where $\varepsilon$ is the complex dielectric function and $\mathbf{D}(\omega)$ is the total electric displacement field.
II.2 Quantum response functions
Let us consider charged plasma comprising of $N$ identical quantum particles with charge e-e in a cubic box of volume $V=L^{3}$. For convenience, and without loss of generality, we can reduce the rank-2 susceptibility tensor $\chi(\omega)$ to its spherically symmetric diagonal, denoted $z$, and consider the response of this (homogeneous, isotropic) system to a long-wavelength electric field $E_{z}(\omega)$ applied along the direction $z$. From the zero-field limit of Eq. (8), we can derive